OpenStudy (anonymous):

how would you calculate the molecular weight of a compound if 4.00g of it plus 50.0 grams of water give a solution with a boiling point of 100.41 C the molal boiling constant for water is 0.52 C x kg.mol ? how would you do this ? IM STUCKKK HELPPPP

OpenStudy (aaronq):

dT=i*m*Kb dT= change in temp i=van't off constant (i guess assume it to be 1 because they don't tell you the compound dissociates) m=molality Kb=molal boiling point constant

OpenStudy (anonymous):

is it like this one ΔTB=KBm @aaronq

OpenStudy (aaronq):

yep its the same thing

OpenStudy (anonymous):

0.52 C x kg.mol x ____ = 100.41?

OpenStudy (aaronq):

no, it's the change in temperature FROM the normal boiling temp, so: (0.52 C kg.mol)*m = 0.41

OpenStudy (aaronq):

from there you can find the moles, then the molecular mass

OpenStudy (anonymous):

0.41?

OpenStudy (aaronq):

dT=NEW BP- NORMAL BP= 100.41-100 = 0.41

OpenStudy (anonymous):

wait why would you take out the 100? @aaronq

OpenStudy (aaronq):

because the formula uses the change (represented by Δ) in boiling point temperature (caused the solute). So you need to use only the change

OpenStudy (anonymous):

how would you know that the change is

OpenStudy (aaronq):

you would find the difference between the new boiling point temp from the normal (without solute) bp. dT=NEW BP- NORMAL BP= 100.41-100 = 0.41

OpenStudy (anonymous):

i just don't understand how to find the normal BP

OpenStudy (aaronq):

the normal boiling point of water is 100 degrees celsius

OpenStudy (anonymous):

so 0.788462 moles/kg os solvent so 0.788462 x (molar mass of H2o? )/mole = 14.2044

OpenStudy (anonymous):

?

OpenStudy (aaronq):

nope, what did you get from the equation above for molality?

OpenStudy (anonymous):

uh isn't molality moles of solute/ kg of solvent so 0.788426m is it not?

OpenStudy (aaronq):

yep thats right, why are you using the molecular mass of water if it's in kg? it should be 0.788426m = n/0.05 kg

OpenStudy (anonymous):

n=0.039421? do you add that to the 4.00 grams?

OpenStudy (aaronq):

nope, the formula for moles is n=m/M so 0.039421=4g/M

OpenStudy (anonymous):

wait what des n m and M stand for?

OpenStudy (aaronq):

n = moles, m = mass, M= molar mass

OpenStudy (anonymous):

so the Molar mass is 0.009855?

OpenStudy (aaronq):

M=4 g/0.039421= 101.4687603054209685 g/mol

OpenStudy (anonymous):

oh whoops THANKS this finally makes sense tytytyty

OpenStudy (aaronq):

good stuff! no problem

OpenStudy (anonymous):

0.788426m = n/0.05 kg why would you set up the equation like this though @aaronq

OpenStudy (anonymous):

WAIT ACTUALLY JK I GOT IT

OpenStudy (aaronq):

haha it just the definition of molality